Question:
Find the seventh term of the arithmetic sequence if the first term is -6 and the common difference is 5.
(A) 29
(B) 24
(C) -36
(D) 34
Thank you!

Answer by venugopalramana(3286)

TN=A+(N-1)D...FOR A.P....WHERE
TN=N TH.TERM
A=FIRST TERM
N=NUMBER OF TERMS
D=COMMON DIFFERENCE....HERE
N=7,A=-6,D=5...
T7=-6+(7-1)5=-6+30=24.......B IS THE ANSWER